OpenStudy (hellokitty17):

HELP ME!!!!!!!! ANYONE WHO IS IN FLVS ALGEBRA 1 JOIN MY QUESTION!!!! I NEED HELP!! PLEASE AND THANK YOU!

OpenStudy (hellokitty17):

The graph below shows a company's profit f(x), in dollars, depending on the price of pens x, in dollars, being sold by the company: graph of quadratic function f of x having x intercepts at ordered pairs 0, 0 and 6, 0. The vertex is at 3, 120 Part A: What do the x-intercepts and maximum value of the graph represent? What are the intervals where the function is increasing and decreasing, and what do they represent about the sale and profit? (5 points) Part B: If at one time the profit of the company was at least $60, what domain could possibly produce this profit? (2 points) Part C: What is an approximate average rate of change of the graph from x = 3 to x = 5, and what does this rate represent? (3 points)

OpenStudy (hellokitty17):

@IrishBoy123

OpenStudy (irishboy123):

diagram / graph?

oregonduck (oregonduck):

Part B: When y is 60 x is both 1 and 5 Part C: We have to look at x = 3 and x = 5 and find the 'y' values..which is 120 and 60 (3, 120) (5, 60) Plug them into the slope formula: m=y2−y1/x2−x1 m=60−120/5−3 Simplify

OpenStudy (hellokitty17):

how to simplify

OpenStudy (hellokitty17):

@IrishBoy123 help me

oregonduck (oregonduck):

the answer is -30

OpenStudy (irishboy123):

so you want help with B and C, right?

OpenStudy (irishboy123):

here's a better graph

OpenStudy (irishboy123):

for B you are looking for a "domain", which here is the set of all the x values that go into your function to give a profit of **at least** $60. it's not 1 and 5.

OpenStudy (irishboy123):

this is how you figure it out

OpenStudy (irishboy123):

it looks like \(a<x<b\) you need to find a,b